OpenStudy (anonymous):

How do you find the expanded form of a binomial? Such as: (5x+9)^2

OpenStudy (gorv):

(5x+9)*(5x+9)=(5x+9)*5x+9*(5x+9)=25x^2+45x+45x+81=25x^2+90x+81

OpenStudy (anonymous):

"how" do you find it

OpenStudy (anonymous):

(5x+ 9)^2 (5x + 9)(5x + 9) use the FOIL method : First, outside, inside, last First : multiply the first terms in each set... 5x * 5x = 25x^2 Outside : multiply the outside terms in each set... 5x * 9 = 45x Inside : multiply the inside terms in each set.. 9 * 5x = 45x Last : multiply the last terms in each set .. 9 * 9 = 81 so now we have : 25x^2 + 45x + 45x + 81 combine like terms and you get : 25x^2 + 90x + 81

OpenStudy (anonymous):

does that help ?

OpenStudy (anonymous):

yes! thank you soooo much! :)

OpenStudy (anonymous):

no problem :)